The magnetic field due to the earth has a horizontal component of 26 μT at a place where the dip is 60°. Find the vertical component and the magnitude of the field.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. δ(dip) = 60° B base H = B cos 60° ⇒ B = 52 * 10^-6 = 52 μT B base v = B sin δ = 52 * 10^-6 √3/2 = 44.98 μT = 45 μT